Project

General

Profile

NDOS Commissioning Meeting 2012/04/09

Attendance: Jon, Athans, Peter, Andrew, Mat, Gavin, Enrique, Minerba, Nate, Denis, Susan,

Agenda:

  1. Beam and Detector Update
  2. DataCheck Report
  3. Installation Plans for the week
  4. DAQ Update and Plans for the week
  5. DCS Update and Plans for the week
  6. AOB

Beam and Detector Report

  • DAQ ran smoothly through this weekend. This morning Jon stopped the run, did a DSO scan and started a new run. A short time later we stopped collecting data from the MC; Peter noticed that there were people working on the detector. A new run was started, but a few hours later a similar problem was found (two DCMs died during the run)

DataCheck Report:

  • none for this week

Installation plans for this week

  • Mat: work on purge system is ~complete.
  • Jon: we should test the robustness of the new DCMs this week by unplugging FEBs

DAQ Update and Plans for next week

  • Peter: want to update software, sometime over the next couple of days.
  • Jon found that run history has not made it to the DAQ dB since 3/27.
  • Peter: VNC is set up on nova-daq-4, which will be for DAQ experts only. Will be very useful for ~remote control rooms.
  • Gavin is tagging the offline now, so NDOS event display can move to a tagged version.

DCS Update and Plans for next week

  • Athans: a few final tests of the IOC on the NDOS
  • Jon: what's going on with pvArchiver? Athans: Kurt is working on it, will likely have a solution to test this week

AOB

  • None